# Superdiffusion and antidiffusion in an aligned active suspension

### v1

@@ Abstract
We show theoretically that an imposed uniaxial anisotropy leads to new universality classes for the dynamics of active particles suspended in a viscous fluid. In the homogeneous state, their concentration relaxes superdiffusively, stirred by the long-ranged flows generated by its own fluctuations, as confirmed by our numerical simulations. Increasing activity leads to an anisotropic diffusive instability, driven by an active contribution to the particle current proportional to the local curvature of the suspension velocity profile.
@@
Active Matter, that is, living materials and their imitations, is made up of components that continually extract mechanical work from a fuel supply [1, 2]. The dynamical equations governing active materials differ from their passive counterparts through physical fluxes arising from maintained chemical forces [3]. In the cases of greatest interest, these fluxes dominate at large length scales, resulting in dynamical and statistical properties qualitatively different from those of a time-reversal-invariant system with the same spatial symmetries. Much insight has emerged recently [4, 5, 6, 7] from a focus on active currents and stresses built from scalar fields [8] rather than alignment or flocking. In this Letter we study such scalar active matter in a permanently anisotropic momentum-conserving fluid, described, like active [9, 6] or passive [10] model H, by concentration and hydrodynamic velocity fields, but with a preferred axis. We discuss possible experimental realizations, e.g., Fig. 1, towards the end of the paper.
@@
Here are our main results. The self-advection of concentration fluctuations by the long-ranged active flows they generate leads to a dynamic exponent $z=d/2$ for $d\leq 4$: inhomogeneities on a length scale $L$ relax on a timescale $\propto L^{3/2}$ in $d=3$. Our numerical studies of Brownian force dipoles in a Stokesian fluid confirm this asymptotic superdiffusive scaling, and uncover an early-time ballistic dynamics that also emerges from our theory. In addition, we point out the existence of an active contribution to solute currents proportional to the curvature in the profile of the hydrodynamic velocity field (Fig.˜2), that can drive an anisotropic diffusive instability, with a critical point at which the hydrodynamic interaction is strongly relevant. We thus uncover two hitherto unknown universality classes, describing the homogeneous phase and the onset of phase separation of uniaxial active model H.
We now show how these results were obtained. We work with an active suspension, with solute concentration field $c({\bf r},t)$ and joint velocity field ${\bf u}({\bf r},t)$, as functions of position $\bf r$} and time $t$, and a macroscopic anisotropy defined by a distinguished $\hat{\bf z}$ direction, with $\hat{\bf z}\to-\hat{\bf z}$ symmetry and isotropy in the $\perp$ plane transverse to $\hat{\bf z}$. We enforce overall incompressibility, i.e., total density $\rho=$ constant, as appropriate for the slow flows of interest here, so that $\nabla\cdot{\bf u}=0$. The imposed, rather than spontaneous, alignment means that $c$ and $\bf u$} are the only slow variables, obeying the conservation laws $\partial_{t}c=-\nabla\cdot{\bf J}$ for number and $\partial_{t}(\rho{\bf u})=\nabla\cdot{\bm{\sigma}}$ for momentum, with current $\bf J$} and stress tensor $\bm{\sigma}$. To build our equations of motion, we need only ask what contributions to $\bf J$} and $\bm{\sigma}$, constructed from $c$ and $\bf u$} and their gradients, can arise in the presence of anisotropy and activity, that compete with those already accounted for in (isotropic) active model H [9, 6]. A little reflection will show that there are precisely two contributions, at linear order in the fields:
@@
to the stress, linear in the concentration, and
@@
which we term Flow-Induced Motility or FIM, to the current. Galilean invariance, incompressibility, and symmetry under inversion of and rotation about $\hat{\bf z}$ rule out any contribution (apart from $c{\bf u}$) of $\bf u$} to the particle current at lower gradient order than (Superdiffusion and antidiffusion in an aligned active suspension). Eqs. (1) and (Superdiffusion and antidiffusion in an aligned active suspension) lead to the equations of motion
@@
$+(D_{\perp}\nabla_{\perp}^{2}+D_{z}\partial_{z}^{2})c+\nabla\cdot\bf f+$... (3)
@@
$\rho(\partial_{t}+{\bf u}\cdot\nabla){\bf u}=\eta\nabla^{2}{\bf u}-{\nabla\mathcal{P}}-W\hat{\bf z}\partial_{z}c+$... (4)
where $a,b$ are combinations of the $\{a_{i}\},\{b_{i}\}$ in Superdiffusion and antidiffusion in an aligned active suspension, the ellipsis denotes terms arising in [9, 6] or trivially anisotropic variants, $D_{\perp}$ and $D_{z}$ are bare diffusivities and $\eta$ is the shear viscosity, taken for simplicity to be isotropic, and the pressure $\mathcal{P}$ enforces incompressibility. In Superdiffusion and antidiffusion in an aligned active suspension, we have allowed for number-conserving fluctuations in the form of a Gaussian random current $\mathbf{f}(\mathbf{r},t)$, white in space and time, with strengths $_{z}, NN_{\perp}$ respectively along and transverse to $\hat{\bf z}$. In Eq.˜4 $W$ is positive (negative) for extensile or pusher (contractile or puller) swimmers. The presence of active forcing linear in $\nabla c$, impossible in the isotropic theory [6], is the central feature of Eq.˜4. Superdiffusion and antidiffusion in an aligned active suspension and 4 constitute Uniaxial Active Model H. For the remainder of this paper, we will disregard inertia by setting the left-hand side of Eq.˜4 to zero, a valid approximation for describing microbial suspensions. We will then see that the velocity field $\bf u$} qualitatively changes the concentration dynamics through (Superdiffusion and antidiffusion in an aligned active suspension) and (1).
@@
Some remarks are in order at this stage. First, our decision to invoke FIM stands vindicated as its effect enters Superdiffusion and antidiffusion in an aligned active suspension at the same order in wavenumber as the bare diffusive dynamics. Further, the diffusivity shifts $aW/\eta$, $bW/\eta$ are quadratic in activity, relying on active contributions to both Superdiffusion and antidiffusion in an aligned active suspension and 4, in contrast to the familiar scenario, see, e.g., [11, 12, 13], of interplay between a passive and an active term.
@@
Symmetry allows either sign for $a,b$, but Stokesian-swimmer physics shows $aW$ and $bW$ are always positive: Fig.˜2 shows flows with (a) $\partial_{\perp}\partial_{z}u_{z}>0$ and (b) $\partial_{z}\partial_{\perp}u_{\perp}>0$. If $a,b>0$, then Superdiffusion and antidiffusion in an aligned active suspension tells us these give rise respectively to currents $J^{u}_{\perp}>0$ (rightward) and $J^{u}_{z}<0$ (downward). Thus, the blobs in (a) and (b) move with their centres of drag ahead and astern respectively. This means (see, e.g., [14]) that with respect to the $\hat{\bf z}$ direction both are pushers, i.e., $W>0$. Similarly, $W<0$ for $a,b<0$. Therefore, aW,bW>0aW,\ bW>0, so that FIM, via Superdiffusion and antidiffusion in an aligned active suspension, always reduces the diffusivity. For large enough $a$ or $b$ the diffusivity will first hit zero for a cone of directions determined by Superdiffusion and antidiffusion in an aligned active suspension. Beyond this threshold $D(\theta)<0$ for a range of $\theta$ – the anti-diffusion alluded to in the title – i.e., a linear instability to small-wavenumber concentration fluctuations. Active hydrodynamics thus leads to phase separation without attractive interactions, with structural anisotropy playing the crucial role of orienting the persistent large-scale flows of the swimmers. The condensation can take place in an infinite system, with no need to invoke boundaries [15]. This discussion of course presumes an appreciable magnitude for $a$ and $b$. In the End Matter, we estimated them to be of order $(W\tau/\eta R)c_{0}$ where $\tau$ is the structural relaxation time of the suspension and $R$ the particle size. See End Matter for a discussion on possible effects of the hydrodynamic interaction at the putative critical point of this flow-induced condensation.
Note that even when the system is confined by walls to a 2D fluid film [16], the active currents generated by the averaged flow field along the confinement direction still contribute diffusion-like terms to the continuity equation Superdiffusion and antidiffusion in an aligned active suspension. This results in the possibility of condensation through the same mechanism [17]. Furthermore, although we assumed instantaneous steady Stokesian dynamics above, introducing viscous retardation for the velocity field – i.e., considering the unsteady Stokes equation – still leads to condensation with precisely the same diffusivity shifts as obtained earlier [17].
@@
To summarise: we have shown that the large-scale, long-time properties of a suspension of active particles in a viscous fluid with a built-in preferred axis lie in universality classes distinct from those of an isotropic active suspension with the same slow variables. Unlike in [6], activity enters even in the linearized dynamics of the concentration and hydrodynamic velocity. In the homogeneous phase, we predict superdiffusive relaxation of the concentration, with dynamic exponent $z=d/2$ in dimension $d<4$, i.e., time $\propto$ (length)3/2 in 3D. Our 3D numerical experiments with aligned Brownian force dipoles in a viscous fluid show a crossover from an early-time ballistic to the asymptotic superdiffusive behavior, confirming our proposed scaling. We have identified an active-particle current induced by curvature in the suspension velocity field, not discussed earlier in the literature, that drives phase separation through hydrodynamic flow. The underlying mechanism is fundamentally multi-particle in nature and thus inaccessible to the treatment of [14]. By contrast with the flow-induced phase separation of [15], it does not rely on sample boundaries. Possible physical realizations on which to test our theory include microbes dispersed homogeneously in a stiff nematic [27], or subjected to polarized [28] or directional [29, 30, 31] light fields, or colloidal suspensions in a uniform electric field [32, 33, 34], in which concentration fluctuations are force dipoles.
Acknowledgments— We thank J K Bhattacharjee, R Golestanian, A Maitra and P Popli for valuable discussions. SKS acknowledges computational support from Supercomputer Education and Research Centre, IISc. SR acknowledges support from a JC Bose Fellowship of the ANRF, India, and thanks the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programmes “New statistical physics in living matter: non equilibrium states under adaptive control” and “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, supported by EPSRC Grant No. EP/R014604/1, and for a Rothschild Distinguished Visiting Fellowship. KVK acknowledges support of the Department of Atomic Energy, Government of India, under Project No. RTI4001.

### v2

@@ Abstract
We show theoretically that an imposed uniaxial anisotropy leads to new universality classes for the dynamics of active particles suspended in a viscous fluid. In the homogeneous state, their concentration relaxes superdiffusively, stirred by the long-ranged flows generated by its own fluctuations, as confirmed by our numerical simulations. Increasing activity leads to an anisotropic diffusive instability, driven by the interplay of active stresses with a particle current proportional to the local curvature of the suspension velocity profile.
@@
Active Matter, that is, living materials and their imitations, is made up of components that continually extract mechanical work from a fuel supply [1, 2]. The dynamical equations governing active materials differ from their passive counterparts through physical fluxes arising from maintained chemical forces [3]. In the cases of greatest interest, these fluxes dominate at large length scales, resulting in dynamical and statistical properties qualitatively different from those of a time-reversal-invariant system with the same spatial symmetries. Much insight has emerged recently [4, 5, 6, 7, 8] from a focus on active currents and stresses built from scalar fields [9] rather than alignment or flocking. In this Letter we study such scalar active matter in a permanently anisotropic momentum-conserving fluid, described, like active [10, 6] or passive [11] model H, by concentration and hydrodynamic velocity fields, but with a preferred axis. We discuss possible experimental realizations, e.g., Fig. 1, towards the end of the paper.
@@
Here are our main results. The self-advection of concentration fluctuations by the long-ranged active flows they generate leads to a dynamic exponent $z=d/2$ for $d\leq 4$: inhomogeneities on a length scale $L$ relax on a timescale $\propto L^{3/2}$ in $d=3$. Our numerical studies of Brownian force dipoles in a Stokesian fluid confirm this asymptotic superdiffusive scaling, and uncover an early-time ballistic dynamics that also emerges from our theory. In addition, we show that solute currents proportional to the curvature in the hydrodynamic velocity profile (Fig.˜2) [12, 13], in concert with the active stresses, can drive a diffusive instability, with a critical point at which the hydrodynamic interaction is strongly relevant. We thus uncover two hitherto unknown universality classes, describing the homogeneous phase and the onset of phase separation of uniaxial active model H.
We now show how these results were obtained. We work with an active suspension, with solute concentration field $c({\bf r},t)$ and joint velocity field ${\bf u}({\bf r},t)$, as functions of position $\bf r$} and time $t$, and a macroscopic anisotropy defined by a distinguished $\hat{\bf z}$ direction, with $\hat{\bf z}\to-\hat{\bf z}$ symmetry and isotropy in the $\perp$ plane transverse to $\hat{\bf z}$. We enforce overall incompressibility, i.e., total density $\rho=$ constant, as appropriate for the slow flows of interest here, so that $\nabla\cdot{\bf u}=0$. The imposed, rather than spontaneous, alignment means that $c$ and $\bf u$} are the only slow variables, obeying the conservation laws $\partial_{t}c=-\nabla\cdot{\bf J}$ for number and $\partial_{t}(\rho{\bf u})=\nabla\cdot{\bm{\sigma}}$ for momentum, with current $\bf J$} and stress tensor $\bm{\sigma}$. To build our equations of motion, we need only ask what contributions to $\bf J$} and $\bm{\sigma}$, constructed from $c$ and $\bf u$} and their gradients, can arise in the presence of anisotropy that compete with those already accounted for in (isotropic) active model H [10, 6]. A little reflection will show that there are precisely two contributions, at linear order in fields. One is a uniaxial active stress [14]
@@
which, given the fixed orientation, operates through inhomogeneities in the concentration. The other, permitted in a passive system, is a current
@@
which we will term Flow-Induced Migration (FIM), and can be viewed as an anisotropic variant of the migration due to inhomogeneous gradients discussed in [12, 13]. Anisotropy is crucial: in the isotropic limit, ignoring the $c$-dependence of $a_{i},b_{i}$ in (Superdiffusion and antidiffusion in an aligned active suspension), $\nabla\cdot{\bf J}^{u}\propto\nabla^{2}\nabla\cdot{\bf u}$ which does not contribute for an incompressible system. Galilean invariance, incompressibility, and symmetry under inversion of, and rotation about, $\hat{\bf z}$ rule out contributions (apart from $c{\bf u}$) of $\bf u$} to the particle current at lower gradient order than (Superdiffusion and antidiffusion in an aligned active suspension). Eqs. (1) and (Superdiffusion and antidiffusion in an aligned active suspension) lead to the equations of motion
@@
$+(D_{\perp}\nabla_{\perp}^{2}+D_{z}\partial_{z}^{2})c+\nabla\cdot\bf f+\ldots,$ (3)
@@
$\rho(\partial_{t}+{\bf u}\cdot\nabla){\bf u}=\eta\nabla^{2}{\bf u}-{\nabla\mathcal{P}}-W\hat{\bf z}\partial_{z}c+\ldots,$ (4)
where $a,b$ are combinations of the $\{a_{i}\},\{b_{i}\}$ in Superdiffusion and antidiffusion in an aligned active suspension, the ellipsis denotes irrelevant contributions including the Onsager partners [13] of (Superdiffusion and antidiffusion in an aligned active suspension) in (4), terms arising in [10, 6] or trivially anisotropic variants, $D_{\perp}$ and $D_{z}$ are bare diffusivities and $\eta$ is the shear viscosity, taken for simplicity to be isotropic, and the pressure $\mathcal{P}$ enforces incompressibility. In Superdiffusion and antidiffusion in an aligned active suspension, we have allowed for number-conserving fluctuations in the form of a Gaussian random current $\mathbf{f}(\mathbf{r},t)$, white in space and time, with strengths $_{z}, NN_{\perp}$ respectively along and transverse to $\hat{\bf z}$. In Eq.˜4 $W$ is positive (negative) for extensile or pusher (contractile or puller) swimmers aligned along $\hat{\bf z}$. The presence of active forcing linear in $\nabla c$, impossible in the isotropic theory [6], is the central feature of Eq.˜4. Superdiffusion and antidiffusion in an aligned active suspension and 4 constitute Uniaxial Active Model H. For the remainder of this paper, we will disregard inertia by setting the left-hand side of Eq.˜4 to zero, a valid approximation for describing microbial suspensions. We will then see that the velocity field $\bf u$} qualitatively changes the concentration dynamics through (Superdiffusion and antidiffusion in an aligned active suspension) and (1).
@@
The inclusion of FIM (Superdiffusion and antidiffusion in an aligned active suspension) stands vindicated as its effect enters Superdiffusion and antidiffusion in an aligned active suspension at the same order in wavenumber as the bare diffusive dynamics.
@@
In addition to the passive contribution from [12, 13], the physics of Stokesian swimmers leads to an active piece of $a$ and $b$ with the same sign as $W$, i.e., a positive contribution to $aW$ and $bW$: Fig.˜2 shows flows with (a) $\partial_{\perp}\partial_{z}u_{z}>0$ and (b) $\partial_{z}\partial_{\perp}u_{\perp}>0$. If $a,b>0$, then Superdiffusion and antidiffusion in an aligned active suspension tells us these give rise respectively to currents $J^{u}_{\perp}>0$ (rightward) and $J^{u}_{z}<0$ (downward). Thus, the blobs in (a) and (b) move with their centres of drag ahead and astern respectively. This means (see, e.g., [15]) that with respect to the $\hat{\bf z}$ direction both are pushers, i.e., $W>0$. Similarly, $W<0$ for $a,b<0$. so that the active piece of FIM, via Superdiffusion and antidiffusion in an aligned active suspension, always reduces the diffusivity. For large enough $aW$ or $bW$ the diffusivity will first hit zero for a cone of directions determined by Superdiffusion and antidiffusion in an aligned active suspension. Beyond this threshold $D(\theta)<0$ for a range of $\theta$ – the anti-diffusion alluded to in the title – i.e., a linear instability to small-wavenumber concentration fluctuations. Active hydrodynamics thus leads to phase separation without attractive interactions, with structural anisotropy playing the crucial role of orienting the persistent large-scale flows of the swimmers. The condensation can take place in an infinite system, with no need to invoke boundaries [16]. This discussion of course presumes an appreciable magnitude for $a$ and $b$. The migration velocity estimate in Ref. [12] would imply $a\sim-b\sim c_{0}R^{2}$, which means $a$ and $b$ respectively promote instability for extensile and contractile systems. In End Matter, we argue that the additional active contribution discussed above should be of order $(W\tau/\eta R)c_{0}$ where $\tau$ is the structural relaxation time of the suspension and $R$ the particle size. We show that the combined relative shift in diffusivities should be of order $\phi Pe[1+\mathcal{O}(Pe)]$, where $\phi$ is the volume fraction and $Pe=W/\eta D_{0}R$ a Péclet number, which we estimate to be of order unity for plausible parameter values. See End Matter also for a discussion on possible effects of the hydrodynamic interaction at the putative critical point of this flow-induced condensation.
@@
To summarise: we have shown that the large-scale, long-time properties of a suspension of active particles in a viscous fluid with a built-in preferred axis lie in universality classes distinct from those of an isotropic active suspension with the same slow variables. Unlike in [6], activity enters even in the linearized dynamics of the concentration and hydrodynamic velocity. In the homogeneous phase, we predict superdiffusive relaxation of the concentration, with dynamic exponent $z=d/2$ in dimension $d<4$, i.e., time $\propto$ (length)3/2 in 3D. Our 3D numerical experiments with aligned Brownian force dipoles in a viscous fluid show a crossover from an early-time ballistic to the asymptotic superdiffusive behavior, confirming our proposed scaling. We have shown that the interplay of active stresses and a particle current induced by inhomogeneous gradients in the suspension velocity field drives phase separation through hydrodynamic flow. In End Matter, we argue that the relative shift in diffusivities is of order the Péclet number associated with activity, which for reasonable parameter values is found to be of order unity. The underlying mechanism is fundamentally multi-particle in nature and thus inaccessible to the treatment of [15]. By contrast with the flow-induced phase separation of [16], it does not rely on sample boundaries. Possible physical realizations on which to test our theory include microbes dispersed homogeneously in a nematic [27] stiff enough that its orientational instability due to active stress [14] is banished to scales beyond the system size—, or subjected to polarized [28] or directional [29, 30, 31] light fields, or colloidal suspensions in a uniform electric field [32, 33, 34], in which concentration fluctuations are force dipoles.
Acknowledgments— We thank J K Bhattacharjee, A Callan-Jones, R Golestanian, F Jülicher, A Maitra and P Popli for valuable discussions. SKS acknowledges computational support from Supercomputer Education and Research Centre, IISc. SR acknowledges support from a JC Bose Fellowship of the ANRF, India, and the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the programmes “New statistical physics in living matter: non equilibrium states under adaptive control” and “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, supported by EPSRC Grant No. EP/R014604/1, and for a Rothschild Distinguished Visiting Fellowship. SR also acknowledges the support and hospitality of the workshop “From Soft Matter to Biophysics”, Les Houches (2023). KVK acknowledges support of the Department of Atomic Energy, Government of India, under Project No. RTI4001.